A weather balloon has a volume of 765L when filled with helium at 10°C added pressure of 395 tour what is the new volume of the balloon in liters if the balloon rises to a point where the air pressure is 235 tour and the temperature is -24°C
Answer #1

According to combined gas law. - Ti P, initial pressure vt initial volume T → initial temperature in Kelvin Py + final pressuP = zastov Pg = 23story V₂ = ? To = 273-24 = 249k. v = 7652 Ti = 10+273=2836 (39stor) (7652) (235 torr) (V₂) (249K) (283K) 39